# RGB2Point: 3D Point Cloud Generation from Single RGB Images

### v1

@@ Abstract
We introduce RGB2Point, an unposed single-view RGB image to a 3D point cloud generation based on Transformer. RGB2Point takes an input image of an object and generates a dense 3D point cloud. Contrary to prior works based on CNN layers and diffusion-denoising approaches, we use pre-trained Transformer layers that are fast and generate high-quality point clouds with consistent quality over available categories. Our generated point clouds demonstrate high quality on a real-world dataset, as evidenced by improved Chamfer distance (51.15%) and Earth Mover’s distance (45.96%) metrics compared to the current state-of-the-art. Additionally, our approach shows a better quality on a synthetic dataset, achieving better Chamfer distance (39.26%), Earth Mover’s distance (26.95%), and F-score (47.16%). Moreover, our method produces 63.1% more consistent high-quality results across various object categories compared to prior works. Furthermore, RGB2Point is computationally efficient, requiring only 2.3GB of VRAM to reconstruct a 3D point cloud from a single RGB image, and our implementation generates the results 15,133$\times$ faster than a SOTA diffusion-based model.
@@ 1 Introduction
We analyze the impact of pre-trained weights on the effectiveness of the ViT for 3D point cloud reconstruction tasks. Moreover, we swap the image feature extractor part to a pre-trained ResNet50 [15] to evaluate the effect of ViT. Our method shows reconstruction improvements, including quantitative and qualitative results over existing works, as RGB2Point outperforms in Chamfer distance (39.26% and 51.15%) and Earth Mover’s distance (26.95% and 45.96%) compared to the current state-of-the-art. Also, our model shows 47.1% higher reconstruction quality in F-score compared to the diffusion-based method [31]. Contrary to previous work, our reconstructed point clouds show a 63.1% more consistently high quality for all object classes. We claim the following contributions: 1) providing a new model architecture that is VRAM efficient but also generates a high-quality point cloud from a single RGB image, 2) a Transformer model can generate higher quality 3D objects than a probabilistic denoising diffusion model. We will share the code and model upon acceptance.
@@ 2 Related Work
Transformer uses the attention [49] mechanism and it has shown its outstanding performance in Natural Language Processing (NLP) such as question answering [9, 61, 59], text generation [3, 47] and sentiment analysis [25, 60]. Transformer [49] has also shown superior performances in the Computer Vision area as it outperforms image classification tasks [10, 17], object detection [4, 12], semantic segmentation [26, 54]. Moreover, the performance of Transformer [49] continues in 3D space as well in different domains such as 3D point cloud completion task [65, 62], 3D reconstruction [51] and 3D tree generation [22]. We utilize the Vision Transformer as an image feature extractor of our model to reconstruct 3D point cloud data. Our approach exhibits improvements across various metrics, including 39.26% and 51.15% for Chamfer Distance, as well as 26.95% and 45.96% for Earth Mover’s distance, evaluated on synthetic object datasets [6] and real object datasets [44]. In contrast to the existing methods [33, 32, 18] that rely on CNN-based feature extractors, our model demonstrates a better performance in 3D generation metrics as we shown in Sec. 4.4 from the synthetic dataset [6] and Sec. 4.4 from the real-world dataset [44].
Diffusion models are new proposing approaches for a generative deep learning model by iterative denoising process. Using denoising diffusion probabilistic models (DDPM) [16] or from a latent space [39]. By utilizing these two fundamental models, they contribute numerous applications such as a novel view synthesis [5, 66] and a 3D object generation [24, 36].
In this emergence of diffusion-based models, researchers leverage its approach to point cloud such as an unconditional diffusion model to generate point clouds [28, 34], point cloud completion using a diffusion model [29] and a single image to point cloud reconstruction [31]. The diffusion methods leverage probabilistic approaches that show a high quality of the unseen field. When the number of reconstructing categories is diverse, one of the critical elements of the model is stability over the categories. From Tab. 4.4, we infer the stability of each model by calculating a standard deviation of F-scores over the reconstructed categories, and the lower value gives a more stable generation quality. Also, we show 15,133 times faster generation along with better qualitative results by comparing a diffusion-based work [34] in Fig. 3.
@@ 3 Approach
Architecture: RGB2Point takes a $224\times 224$ RGB image as its input and generates its corresponding 3D point cloud with $N$ points. From the evaluation, Sect. 4.4, we show that our model is flexible on the number of output point clouds, such as 1024 or 8192 points. We do not need any additional set of layers but simply change the size of the output layer during the training. RGB2Point consists of three parts. The first part is composed of a pre-trained vision Transformer [10] that extracts image features. Next, it has Contextual Feature Integrator, which uses a feed-forward layer with size $A$, and a multi-head attention mechanism with $H$ heads. This module provides the weight of specific regions of weight for generation in each category. The last part is the Geometric Projection Module, which handles 3D mapping from weighted features to 3D space. Especially when we use explicit point cloud space. It has three layers of the feed-forward layer with $D$ dimension with Leaky ReLU [30] activation function with a slope value of 0.2. We set $A=1024$, $H=4$, and $D=2048$. We choose these values by conducting extensive experiments as we show in Sec. 4.5.
@@ 4.4 Evaluation
We validated RGB2Point using two datasets: a synthetic dataset ShapeNet [6] and real-world images from Pix3D [44]. We visualize the generated results of a single RGB image from ShapeNet [6] in Fig. 2 for a qualitative evaluation. However, the qualitative results with a side-by-side comparison with other works were not possible because of the outdated software (Python 2 version) and hardware support (CUDA 8 version) that our workstation cannot run. We contacted authors to obtain pre-trained models but could not get them from them. However, the authors provide their datasets so we can evaluate them quantitatively. Also, we compare qualitative results between a diffusion-based work [34] in complex real-world data in Fig. 3. In the test stage, RGB2Point demonstrates efficiency, utilizing merely 2.34 GB of VRAM when operating with a batch size of one. Furthermore, a diffusion-based model [34] takes 37 minutes and 50 seconds per image on average, but RGB2Point only takes 0.15 seconds per image on average. In other words, our model is 15,133 faster but generates a higher-quality point cloud in a complex real-world single image, as we show in Fig. 3.
@@
First, we compare our generations to a Diffusion-based approach [31] and prior works that generate a 3D occupancy grid, voxel, from a single-view image [7, 58, 56]. All the works are evaluated using 8192 points, and we replace the output layer size with 8,192 instead of 1,024. Based on the F-score that we show in Sec. 4.4, our model shows 47.2% better than the current state-of-the-art model [31] that based on a probabilistic diffusion model which is recently showing high-quality 3D object generation. The SOTA method generates high-quality point clouds on specific categories, such as airplanes or rifles, but its overall stability over the categories is imbalanced. Unlike the biased performance, RGB2Point generates 64.1% more stable generation quality than the SOTA method [31]. We calculate the performance stability using a standard deviation (Stdev.) and report it under the last row in Sec. 4.4. This shows that our model is robust and capable of generating a high-quality, dense point cloud even if we expand the size of the output layer.
@@
Figure 3: Generated point cloud data by RGB2Point using images from the real-world dataset Pix3D [44]. The first column shows an input RGB image, and the second left column shows a generated point cloud from Point-E [34]. The third left column shows generated point cloud data by RGB2Point and the rightmost column shows its ground truth point cloud data.
@@
Table 3: We test our model on the real 3D dataset Pix3D [44], and compared to the related works. The best values are shown in bold and the current SOTA is underlined. RGB2Point shows an improvement of 51.15% in Chamfer distance (CD) and 45.96% in Earth Mover’s distance, compared to the previous state-of-the-art metrics.
@@ 4.5 Ablation Study
We conduct three additional experiments: 1) evaluating the influence of various parameter configurations, 2) assessing how the presence of the pre-trained ViT weights affects the performance of 3D generation, and 3) validating the effectiveness of ViT for point cloud generation task by replacing it to ResNet-50 [15].
@@ 4.6 Limitation
Figure 4: Three failure cases from a complex real-world dataset, [44], with their input images, a generated 3D point cloud from a diffusion-based method [34], ours, and the ground truth 3D point cloud data.
We report failure cases using images from a real-world dataset [44] in Fig. 4. The common issue (identified inFig. 4 as (1-3)) is the lack of level of detail in our generated point cloud data. For (1), RGB2Point generated a chair with shorter legs than the ground truth due to the limitation of viewpoint, especially a top view in this case. Some parts of the back support are missing in the cases of (2) and (3). Compared to our failure generations, a diffusion-based method [34] cannot generate a chair at all from this real-world case scenario.
@@ 5 Conclusions and Future Works
We introduce a fast, high-quality 3D point cloud generative model from a single image. Leveraging the synthetic dataset from Shapenet [6], RGB2Point surpassed both Chamfer distance and Earth Mover’s distance by 39.26% and 26.95% respectively. In addition, our work shows an improvement on the real-world dataset [44] outperforming the current SOTA by 51.15% and 45.96% using Chamfer distance and Earth Mover’s distance. In addition to the higher quality generated point cloud, our model shows a 15,133 times faster inference time than eye-catching diffusion-based models [34, 31].
@@
Additionally, our model can be extended to handle multi-view images for enhanced 3D point cloud generation. A potential improvement involves integrating a differentiable renderer into RGB2Point to apply RGB texture to the point cloud, enhancing visual fidelity. Given its VRAM efficiency, future work includes adapting this AR or VR device deployment model, enabling on-device usage. Furthermore, RGB2Point achieves point cloud generation in approximately 0.15 seconds per image. Considering this efficiency, the model could be quantized and optimized for edge devices, facilitating real-time single-image conditioned generation. Such capabilities could benefit applied robotics, particularly path planning and object evasion maneuvers.

### v3

@@ Abstract
We introduce RGB2Point, an unposed single-view RGB image to a 3D point cloud generation based on Transformer. RGB2Point takes an input image of an object and generates a dense 3D point cloud. Contrary to prior works based on CNN layers and diffusion-denoising approaches, we use pre-trained Transformer layers that are fast and generate high-quality point clouds with consistent quality over available categories. Our generated point clouds demonstrate high quality on a real-world dataset, as evidenced by improved Chamfer distance (51.15%) and Earth Mover’s distance (36.17%) metrics compared to the current state-of-the-art. Additionally, our approach shows a better quality on a synthetic dataset, achieving better Chamfer distance (39.26%), Earth Mover’s distance (26.95%), and F-score (47.16%). Moreover, our method produces 63.1% more consistent high-quality results across various object categories compared to prior works. Furthermore, RGB2Point is computationally efficient, requiring only 2.3GB of VRAM to reconstruct a 3D point cloud from a single RGB image, and our implementation generates the results 15,133$\times$ faster than a SOTA diffusion-based model.
@@ 1 Introduction
We analyze the impact of pre-trained weights on the effectiveness of the ViT for 3D point cloud reconstruction tasks. Moreover, we swap the image feature extractor part to a pre-trained ResNet50 [15] to evaluate the effect of ViT. Our method shows reconstruction improvements, including quantitative and qualitative results over existing works, as RGB2Point outperforms in Chamfer distance (39.26% and 51.15%) and Earth Mover’s distance (26.95% and 36.17%) compared to the current state-of-the-art. Also, our model shows 47.1% higher reconstruction quality in F-score compared to the diffusion-based method [33]. Contrary to previous work, our reconstructed point clouds show a 63.1% more consistently high quality for all object classes. We claim the following contributions: 1) providing a new model architecture that is VRAM efficient but also generates a high-quality point cloud from a single RGB image, 2) a Transformer model can generate higher quality 3D objects than a probabilistic denoising diffusion model. We will share the code and model upon acceptance.
@@ 2 Related Work
Transformer uses the attention [53] mechanism and it has shown its outstanding performance in Natural Language Processing (NLP) such as question answering [9, 65, 63], text generation [3, 50] and sentiment analysis [27, 64]. Transformer [53] has also shown superior performances in the Computer Vision area as it outperforms image classification tasks [10, 18], object detection [4, 12], semantic segmentation [28, 58]. Moreover, the performance of Transformer [53] continues in 3D space as well in different domains such as 3D point cloud completion task [69, 66], 3D reconstruction [55] and 3D tree generation [24]. We utilize the Vision Transformer as an image feature extractor of our model to reconstruct 3D point cloud data. Our approach exhibits improvements across various metrics, including 39.26% and 51.15% for Chamfer Distance, as well as 26.95% and 36.17% for Earth Mover’s distance, evaluated on synthetic object datasets [6] and real object datasets [46]. In contrast to the existing methods [35, 34, 19] that rely on CNN-based feature extractors, our model demonstrates a better performance in 3D generation metrics as we shown in Sec. 4.4 from the synthetic dataset [6] and Sec. 4.4 from the real-world dataset [46].
Diffusion models are new proposing approaches for a generative deep learning model by iterative denoising process. Using denoising diffusion probabilistic models (DDPM) [16] or from a latent space [41]. By utilizing these two fundamental models, they contribute numerous applications such as a novel view synthesis [5, 70] and a 3D object generation [26, 38, 23, 17, 49].
In this emergence of diffusion-based models, researchers leverage its approach to point cloud such as an unconditional diffusion model to generate point clouds [30, 36, 51], point cloud completion using a diffusion model [31] and a single image to point cloud reconstruction [33]. The diffusion methods leverage probabilistic approaches that show a high quality of the unseen field. When the number of reconstructing categories is diverse, one of the critical elements of the model is stability over the categories. From Tab. 4.4, we infer the stability of each model by calculating a standard deviation of F-scores over the reconstructed categories, and the lower value gives a more stable generation quality. Also, we show 15,133 times faster generation along with better qualitative results by comparing a diffusion-based work [36] in Fig. 3.
@@ 3 Approach
Architecture: RGB2Point takes a $224\times 224$ RGB image as its input and generates its corresponding 3D point cloud with $N$ points. From the evaluation, Sect. 4.4, we show that our model is flexible on the number of output point clouds, such as 256, 1024, or 8192 points. We do not need any additional set of layers but simply change the size of the output layer during the training. RGB2Point consists of three parts. The first part is composed of a pre-trained vision Transformer [10] that extracts image features.
Contextual Feature Integrator (CFI) is a designed module to enhance the representation of specific regions within an image, which is important for generating accurate outputs. It consists of a feed-forward layer with size $A$ and a multi-head attention mechanism with $H$ heads. The feed-forward layer acts as a transformation network, mapping the input features into a higher-dimensional space that captures relationships among the features. The dimensionality, $A$, allows the network to model complex dependencies that might not be apparent in the original feature space. By applying non-linear activation functions, this layer introduces non-linearities into the model to learn complex patterns. The multi-head attention mechanism in the model is composed of $H$ heads, each responsible for attending to different aspects of the input feature sequences. This method enables the model to capture a wide range of contextual information within the input sequence, allowing it to attend to multiple relevant features concurrently. Each attention head computes a weighted sum of the input features, where the weights are derived from the similarity between the feature vectors, which serve as both queries, keys, and values in a self-attention setup. The outputs from all attention heads are concatenated and passed through a linear layer, which projects them into a dimensional space appropriate for point cloud generation. This step ensures that the model effectively integrates and highlights important features, resulting in a more precise representation in the generated point cloud.
Geometric Projection Module (GPM) plays a key component in bridging the transition from high-dimensional feature representations to accurate 3D point cloud generation. This module maintains the spatial coherence and structural information of the input features to perform the effective translation of abstract feature spaces into meaningful geometric representations. Following the multi-head attention mechanism, this module employs a sequence of linear layers with Leaky ReLU activations [32] with a slope value of 0.2 to introduce necessary non-linearities and capture complex feature interactions. The module starts with the initial linear transformation, which projects the attention-weighted features into a higher-dimensional space. This step improves the model’s capacity to capture complex spatial relationships and dependencies within the input data. Subsequent linear layers then refine these features, progressively distilling the essential information needed for accurate 3D reconstruction. The final layer of the module performs a critical projection, mapping the refined features into a three-dimensional coordinate space. This ensures that each output point corresponds to a precise $x$, $y$, and $z$ coordinate in the generated point cloud.
We set $A=1,024$, $H=4$, and $D=2,048$. We choose these values by conducting extensive experiments, as shown in Sec. 4.5.
@@ 4.4 Evaluation
We validated RGB2Point using two datasets: a synthetic dataset ShapeNet [6] and real-world images from Pix3D [46]. We visualize the generated results of a single RGB image from ShapeNet [6] in Fig. 2 for a qualitative evaluation. However, the qualitative results with a side-by-side comparison with some other works were not possible because of the outdated software (Python 2 version) and hardware support (CUDA 8 version) that our workstation cannot run. We contacted authors to obtain pre-trained models but could not get them from them. However, the authors provide their datasets so we can evaluate them quantitatively. Also, we compare qualitative results between a recent single image-based 3D reconstruction works [17, 49, 36, 51] in complex real-world data in Fig. 3. In the test stage, RGB2Point demonstrates efficiency, utilizing merely 2.34 GB of VRAM when operating with a batch size of one. Furthermore, a diffusion-based model [36] takes 37 minutes and 50 seconds per image on average, but RGB2Point only takes 0.15 seconds per image on average. In other words, our model is 15,133 faster but generates a higher-quality point cloud in a complex real-world single image, as we show in Fig. 3.
@@
First, we compare our generations to a Diffusion-based approach [33] and prior works that generate a 3D occupancy grid, voxel, from a single-view image [7, 62, 60]. All the works are evaluated using 8192 points, and we replace the output layer size with 8,192 instead of 1,024. Based on the F-score that we show in Sec. 4.4, our model shows 47.2% better than the current state-of-the-art model [33] that based on a probabilistic diffusion model which is recently showing high-quality 3D object generation. The SOTA method generates high-quality point clouds on specific categories, such as airplanes or rifles, but its overall stability over the categories is imbalanced. Unlike the biased performance, RGB2Point generates 64.1% more stable generation quality than the SOTA method [33]. We calculate the performance stability using a standard deviation (Stdev.) and report it under the last row in Sec. 4.4. All the models are trained on the same ShapeNet [6]. This shows that our model is robust and capable of generating a high-quality, dense point cloud even if we expand the size of the output layer.
@@
Figure 3: Generated point cloud data by RGB2Point using images from the real-world dataset Pix3D [46]. The first column shows an input RGB image, and the next two columns show a reconstructed mesh from LRM [17], TripoSR [49]. The third and fourth columns show reconstructed point clouds from Point-E [36] and LION [51]. The sixth left column shows generated point cloud data by RGB2Point and the column with GT shows its ground truth point cloud data. The red arrows highlight differences compared to GT. Also, we show a rotated view from our outputs in the last column.
@@
Table 3: We test our model on the real 3D dataset Pix3D [46], and compared to the related works. The best values are shown in bold and the current SOTA is underlined. RGB2Point shows an improvement of 51.15% in Chamfer distance (CD) and 36.17% in Earth Mover’s distance, compared to the previous state-of-the-art metrics.
@@ 4.5 Ablation Study
We conduct five additional experiments: 1) evaluating the influence of various parameter configurations, 2) assessing how the presence of the pre-trained ViT weights affects the performance of 3D generation, 3) validating the effectiveness of ViT for point cloud generation task by replacing it to ResNet-50 [15], 4) evaluating the effectiveness of two modules: Contextual Feature Integrator and Geometric Projection Module, and 5) qualitative analysis of outputs using different $n$, (the number of points) where $n=$ 128.
@@ H D A Car Aircraft Chair Car Aircraft Chair CD EMD
Moreover, we validate the effectiveness of our two modules by training without them. We evaluate two models that remain the same but remove each module from the entire proposed architecture. We report a significant performance drop compared to metrics from the original architecture in  Sec. 4.5. Even if a module is removed, Sec. 4.4 shows that our approach is at least 39.30% and 9.88% better in Chamfer Distance and Earth Mover’s distance compared to the previous works.
{NiceTabular}
p0.8cm — p1.2cm p1.2cm — p1.2cm p1.2cm
Module   ShapeNet   Pix3D
CD EMD CD EMD
CFI -19.27% -54.43% -21.86% -22.73%
GPM -31.30% -55.06% -26.00% -41.20%
Table 7: We validated the effectiveness of our two modules by removing them from the model architecture, and we show the average of all scenes from this paper. The performances were measured using Chamfer Distance and Earth Mover’s Distances. When we removed the modules CFI and GPM, the performance dropped compared to metrics from the original pipeline.
We conduct an additional experiment by varying the number of point clouds, $n$. Specifically, we reduced the number of points to test whether this decrease affects the preservation of the object’s overall shape during generation. As demonstrated in Fig. 4, even with a smaller $n$, the overall shapes are still generated accurately without losing details.
Figure 4: We compare the output of different numbers of point clouds. Our original pipeline generates 1,024 point clouds but we show 128 point clouds. The overall shape is preserved instead of missing a random region of point clouds.
@@ 4.6 Limitation
Figure 5: Three failure cases from a complex real-world dataset, [46], with their input images, single image-based 3D reconstructions as a mesh [17, 49] and 3D point cloud [36, 51], ours, and the ground truth 3D point cloud data.
We report failure cases using images from a real-world dataset [46] in Fig. 5. The common issue (identified inFig. 5 as (1-4)) is the lack of level of detail in our generated point cloud data. For (1), RGB2Point generated a chair with shorter legs than the ground truth due to the limitation of viewpoint, especially a top view in this case. Some parts of the back support are missing in the cases of (2) and (3). And (4) shows an occlusion that is about 50% of its original shape by a desk. Compared to our failure generations, mesh reconstruction methods [17, 49] do not generate any related objects. Point-E [36] cannot generate a chair at all from this real-world case scenario. LION  [51] gives better quality of point cloud than Point-E  [36] but still it has low accuracy based on the given RGB image as we show this behavior in  Sec. 4.4.
@@ 5 Conclusions and Future Works
We introduce a fast, high-quality 3D point cloud generative model from a single image. Leveraging the synthetic dataset from Shapenet [6], RGB2Point surpassed both Chamfer distance and Earth Mover’s distance by 39.26% and 26.95% respectively. In addition, our work shows an improvement on the real-world dataset [46] outperforming the current SOTA by 51.15% and 36.17% using Chamfer distance and Earth Mover’s distance. In addition to the higher quality generated point cloud, our model shows a 15,133 times faster inference time than eye-catching diffusion-based models [36, 33].
@@
Also, we validate the effectiveness of our modules, Context Feature Integrator and Geometric Projection Module, in Sec. 4.5 that shows an average of 23.93% and 31.97% performance drop using the real datset [46] in Chamfer Distance and Earth Mover’s Distance.
Based on the performance and efficiency of our model, it can be used as a prior or proxy before getting actual lidar-sensor scanned data. Scanning with lidar typically requires multi-view scans or object rotation. Our method generates a high-quality 3D point cloud from a single image in just 0.1 seconds, offering a fast and accurate alternative before full sensor-based scans, which require more time for post-processing. While multi-view approaches can improve performance, they increase memory use and runtime due to model complexity. In contrast, our method needs only 2.3GB of VRAM and 0.1 seconds on a standard GPU, providing an efficient tradeoff.
@@
Additionally, a potential improvement involves integrating a differentiable renderer into RGB2Point to apply RGB texture to the point cloud, enhancing visual fidelity. Given its VRAM efficiency, future work includes adapting this AR or VR device deployment model, enabling on-device usage. Furthermore, RGB2Point achieves point cloud generation in approximately 0.15 seconds per image. Considering this efficiency, the model could be quantized and optimized for edge devices, facilitating real-time single-image conditioned generation. Such capabilities could benefit applied robotics, particularly path planning and object evasion maneuvers.